OpenStudy (anonymous):

Four employees are cleaning out four different wishing wells in an amusement park

OpenStudy (maddy1251):

Is there more to this problem?

OpenStudy (anonymous):

Yes. Sorry. Tracy collects 5.80 more than asuka. Kyle collects half as much as Tracy. Beverly collects 3 times as much as kyle. How much do each collect if the total collected is 246.20

OpenStudy (maddy1251):

That's okay!

OpenStudy (maddy1251):

Word problems are fun. :P Is this Multiple choice or a writing question?

OpenStudy (anonymous):

Writing. The assignment is to define a variable and write algebraic expressions for the given quantities. Then solve

OpenStudy (maddy1251):

Tracy collects 5.80 more than asuka. Kyle collects half as much as Tracy. Beverly collects 3 times as much as kyle. How much do each collect if the total collected is 246.20 Tracy = 5.80+A Kyle = 1/2t Beverly= 3k. Total = 246.20 Does it give Asuka number?

OpenStudy (maddy1251):

Oh well I have the variables and need to set them to 246.20

OpenStudy (maddy1251):

For Beverly, e.g.; 3k=246.20 Divide each side by 3 to get about 82.07. So, Beverly collected $82.07 roughly

OpenStudy (maddy1251):

Kyle = 1/2t 1/2t=246.20 \[\frac{ 1 }{ 2 }t=246.20\]

OpenStudy (maddy1251):

Kyle got $123.10.

OpenStudy (maddy1251):

Oh I messed up. No wonder my numbers were huge.

OpenStudy (maddy1251):

Anyways, start over. Asuka is variable 'x'

OpenStudy (maddy1251):

Tracy is x+5.80 (I was right there)

OpenStudy (maddy1251):

Kyle is 1/2(x+5.80)

OpenStudy (maddy1251):

Beverly would be 3[1/2(x+5.80)]

OpenStudy (maddy1251):

Tracy = $63. Asoka = $57.20 Kyle = $31.50 Bev = $94.50. Right?

OpenStudy (anonymous):

Do you set that to 246.20 then solve?

OpenStudy (maddy1251):

Yup yup

OpenStudy (anonymous):

Do you divide the equation by 2 to get rid of the 1/2 fraction

OpenStudy (anonymous):

Can you show me how you get that?